### v1

# Copositive criteria for a two-component dark matter model

Abstract
We consider a two-component scalar dark matter model in this work, where the scalars are stabilized by extra $Z_{2}\times Z^{\prime}_{2}$ symmetry. To guarantee the stability of the vacuum, we consider the copositive criteria and different choices of the signs of the couplings will contribute to different viable parameter spaces. Based on the copositive criteria, we systematically analyze the possible conditions and pick up 17 different cases with fixing some parameters to be negative. We randomly scan the parameter space under dark matter relic density constraint and direct detection constraint and focus on the quartic couplings with $\lambda_{13}$,$\lambda_{23}$,$\lambda_{14}$ and $\lambda_{24}$. The shapes of the viable parameter space for $|\lambda_{14}|-|\lambda_{24}|$ are almost similar among these cases, while for $|\lambda_{13}|$ and $|\lambda_{23}|$, the larger value are excluded as long as $\lambda_{13}\leqslant 0$ and $\lambda_{23}\leqslant 0$ due to the large interference effect of the $2\to 2$ processes.
1 Introduction
The Standard Model (SM) of particle physics has been remarkably successful in describing the fundamental particles and their interactions. However, the astrophysical observations indicate the existence of dark matter, which constitutes approximately $26.5\%$ of the universe’s mass-energy content Planck:2015fie . The existence of dark matter can be inferred from its gravitational effects on the universe’s large-scale structure, yet its particle nature remains elusive. From the view of particle physics, dark matter can be assumed to be some stable particle carrying no SM gauge group quantum number.
One of the most popular and attractive dark matter candidates is weakly interacting massive particles(WIMPs)Qi:2021rhh ; Qi:2022fzs ; Qi:2021rpa , where the observed dark matter relic density is generated via the so-called “Freeze-out” mechanismRoszkowski:2017nbc . Among the different WIMP scenarios, the singlet scalar dark matter model is the simplest where a new singlet scalar as dark matter is introduced to the SM. The singlet scalar is stabilized by extra $Z_{2}$ symmetry, and there are only two free model parameters in the model with dark matter mass and the quartic coupling of dark matter with SM Higgs. According to Higgs invisible decay search at LHCATLAS:2016neq , dark matter relic density, direct detection constraint He:2016mls ; Escudero:2016gzx and indirect detection constraint arising from gamma-ray spectrum induced by DM annihilation at Fermi-LAT Fermi-LAT:2015kyq ; Fermi-LAT:2015att and HESSHESS:2013rld , the viable dark matter mass region is limited stringently with resonant mass being about half of the SM Higgs mass and high mass region larger than about 3 TeV Wu:2016mbe . The current direct detection experiments have not yielded a clear signal, imposing stringent constraints on DM-nucleon interactions, so the WIMP scenarios are facing a serious crisis.
One possible solution to the problem is the multi-component dark matter models, which include two or more dark matter species. In the multi-component dark matter models, it is possible that one of the dark matter candidates has a large elastic cross-section on nuclei but forms only a subdominant DM component leading to a suppressed signal Belanger:2022qxt therefore escapes from the direct detection constraint. There is no preference to assume that there is only one dark matter candidate in the universe, and discussion about multi-component dark matter can be found in Coleppa:2023vfh ; Zurek:2008qg ; Costa:2022lpy and so on. Particularly, for the multi-component scalar dark matter model, one can introduce extra symmetry such as $Z_{2}\times Z_{2}^{\prime}$hattacharya:2019fgs , $Z_{5}$ symmetry Belanger:2020hyh ; Qi:2024uiz ; Qi:2023egb ,$Z_{7}$ symmetry Belanger:2022esk or other $Z_{N}$ symmetry Bhattacharya:2024nla ; Yaguna:2019cvp ; Yaguna:2021vhb to ensure the stability of dark matter particles.
When constructing a viable multi-component scalar DM model, the stability of the electroweak (EW) vacuum should be guaranteed. For the scalar field theories, the concept of vacuum stability is intricately linked to the properties of the scalar potential, particularly its quartic couplings. A scalar potential that is bounded from below is a prerequisite for a stable vacuum, and this is where the copositive criteria come into play. Copositive matrices, a class of matrices that are positive on non-negative vectors, provide a mathematical framework to assess the stability conditions of scalar potentials. This approach allows for the derivation of analytic necessary and sufficient conditions for vacuum stability, which is essential for ensuring that the scalar potential does not lead to unphysical situations such as a vacuum that is not the global minimum of the potential.
Discussion about copositive criteria and dark matter can be found in DuttaBanik:2020jrj ; Kannike:2012pe ; Chakrabortty:2013mha ; Arhrib:2011uy , and in this work, we consider a simple two-component scalar dark matter model and discuss the copositive criteria on the model. We introduce two singlet scalars $S_{1}$ and $S_{2}$ as dark matter to the SM, which are stabilized by $Z_{2}\times Z_{2}^{\prime}$ symmetry. In addition, another singlet scalar $S_{3}$ with non-zero vacuum expectation value is also introduced to the model, therefore, a $4\times 4$ copositive matrix of the quartic couplings can be constructed in the model. Note that the copositive criteria of the rank $\leqslant 3$ matrix can be given concretely, while for the higher rank, related discussions about the models are always ignored for their complexity. In this paper, we analyze the copositive criteria of the model systematically. Generally speaking, there are eight cases of the copositive criteria for the rank 4 matrix, and one can have more complex cases when considering the signs of the couplings of the model. We pick up 17 different cases from the point of exchange symmetry for simplicity and estimate the viable parameter space within the dark matter relic density constraint and direct detection constraint. The signs of the couplings will not only determine the copositive criteria but also make a difference in the cross section of $2\to 2$ processes related to the dark matter annihilation into other scalars due to the interference effect. Within the copositive criteria, we randomly scan the chosen parameter space and focus on the quartic couplings with $\lambda_{13}$,$\lambda_{23}$,$\lambda_{14}$ and $\lambda_{24}$. We found the shapes of the viable parameter space for $|\lambda_{14}|-|\lambda_{24}|$ are almost similar among these cases, while for $|\lambda_{13}|$ and $|\lambda_{23}|$, the larger value are excluded as long as $\lambda_{13}\leqslant 0$ and $\lambda_{23}\leqslant 0$ due to the large interference effect.
The paper is organized as follows. We give the two-component scalar dark matter in Sec.2. In Sec.3, we discuss the copositivity of the symmetric matrix with eight different cases. In Sec.4, we consider the dark matter phenomenology of the model including relic density and direct detection. In Sec.5, we estimate the parameter space from the view of copositive criteria, we pick up 17 different permutations of the indexes and give the viable parameter space satisfying relic density and direct detection constraints. Finally, we give a summary in Sec.6.
2 Model description
In this section, we consider a simple two-component scalar dark matter model with $Z_{2}\times Z_{2}^{\prime}$ symmetry by introducing three singlet scalars $S_{1,2,3}$, where the corresponding charges for dark matter particles $S_{1}$ and $S_{2}$ are $(-1,1)$ and $(1,-1)$ and $S_{3}$ has non-zero vacuum expectation value (vev), while the SM particles carry $(1,1)$ charge. In addition, we consider another $Z_{2}^{\prime\prime}$ symmetry for $S_{3}$ for simplicity, which can arise from a spontaneously broken $U(1)$ gauge symmetry. Therefore, the potential part of the scalars can be given by:
$\mathcal{V}$ $=\frac{1}{2}M_{1}^{2}S_{1}^{2}+\frac{1}{2}M_{2}^{2}S_{2}^{2}-\mu_{3}^{2}|S_{3}|^{2}-\mu_{H}^{2}|H|^{2}+\lambda_{H}|H|^{4}+\frac{1}{4}\lambda_{11}S_{1}^{4}+\frac{1}{4}\lambda_{22}S_{2}^{4}+$
$\lambda_{33}|S_{3}|^{4}+\lambda_{12}S_{1}^{2}S_{2}^{2}+\lambda_{13}S_{1}^{2}|S_{3}|^{2}+\lambda_{14}S_{1}^{2}|H|^{2}+\lambda_{23}S_{2}^{2}|S_{3}|^{2}+\lambda_{24}S_{2}^{2}|H|^{2}$
$+\lambda_{34}|S_{3}|^{2}|H|^{2}$ (1)
where $H$ is the SM Higgs doublet. Under unitarity gauge, $H$ and $S_{3}$ can be expressed with:
$H=\left(\begin{array}[]{c}0\\ \frac{v_{0}+h}{\sqrt{2}}\end{array}\right)\,,\quad S_{3}=s_{3}+v_{1}\,,\quad$ (2)
where $v_{0}=246$ GeV corresponds to the electroweak symmetry breaking vev and $v_{1}$ is the vev of $S_{3}$. After spontaneous symmetry breaking (SSB), the masses of $S_{1}$ and $S_{2}$ can be given by:
$m_{1}^{2}=M_{1}^{2}+2\lambda_{13}v_{1}^{2}+\lambda_{14}v_{0}^{2},m_{2}^{2}=M_{2}^{2}+2\lambda_{23}v_{1}^{2}+\lambda_{24}v_{0}^{2}$ (3)
where $m_{1}(m_{2})$ represents the mass of $S_{1}(S_{2})$. On the other hand, we have the squared mass matrix of $s_{3}$ and $h$ with:
$\mathcal{M}=\left(\begin{array}[]{cc}2\lambda_{33}v_{1}^{2}&\lambda_{34}v_{0}v_{1}\\ \lambda_{34}v_{0}v_{1}&2\lambda_{H}v_{0}^{2}\\ \end{array}\right).$ (6)
The physical masses of the two Higgs states $h_{1},h_{2}$ are then
$m^{2}_{h_{1}}$ $=\lambda_{H}v_{0}^{2}+\lambda_{33}v_{1}^{2}-\sqrt{(\lambda_{H}v_{0}^{2}-\lambda_{33}v_{1}^{2})^{2}+(\lambda_{34}v_{0}v_{1})^{2}},$
$m^{2}_{h_{2}}$ $=\lambda_{H}v_{0}^{2}+\lambda_{33}v_{1}^{2}+\sqrt{(\lambda_{H}v_{0}^{2}-\lambda_{33}v_{1}^{2})^{2}+(\lambda_{34}v_{0}v_{1})^{2}}$ (7)
The mass eigenstate ($h_{1},h_{2})$ and the gauge eigenstate ($h,s_{3}$) can be related via
$\begin{pmatrix}h_{1}\\ h_{2}\end{pmatrix}=\begin{pmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{pmatrix}\begin{pmatrix}h\\ s_{3}\end{pmatrix}.$ (8)
where
$\tan 2\theta=\frac{\lambda_{34}v_{0}v_{1}}{\lambda_{33}v_{1}^{2}-\lambda_{H}v_{0}^{2}}$ (9)
Furthermore, we can assume $h_{1}$ is the observed SM Higgs and $h_{2}$ is the new Higgs in our model. One can choose the masses of the Higgs particles $m_{h_{1}}$ and $m_{h_{2}}$ as the inputs so that the couplings of $\lambda_{H}$, $\lambda_{33}$ and $\lambda_{34}$ can be given by:
$\lambda_{H}$ $=\frac{(m_{h_{1}}^{2}+m_{h_{2}}^{2})-\cos 2\theta(m_{h_{2}}^{2}-m_{h_{1}}^{2})}{4v_{0}^{2}},$
$\lambda_{33}$ $=\frac{(m_{h_{1}}^{2}+m_{h_{2}}^{2})+\cos 2\theta(m_{h_{2}}^{2}-m_{h_{1}}^{2})}{4v_{1}^{2}},$ (10)
$\lambda_{34}$ $=\frac{\sin 2\theta(m_{h_{2}}^{2}-m_{h_{1}}^{2})}{2v_{0}v_{1}}$
3 Copositivity of symmetric matrix
The extended scalar field models beyond the SM can contain complicated quartic couplings, thus it is important to formulate the criteria that define the boundedness of the potential with the largest parameter space from the view of the vacuum stability constraint. One possible idea is to construct the quartic couplings as a pure square of the combinations of bilinear scalar fields and set their coefficients to be non-negative. Here we have imported the idea of copositivity of symmetric matrices. In our model, the scalar potential quartic terms can be given with a symmetric matrix as follows:
$\mathcal{S}=\left(\begin{array}[]{cccc}\lambda_{11}&\lambda_{12}&\lambda_{13}&\lambda_{14}\\ &\lambda_{22}&\lambda_{23}&\lambda_{24}\\ &&\lambda_{33}&\lambda_{34}\\ &&&\lambda_{44}\end{array}\right).$ (15)
In the following discussion, we use $\lambda_{44}$ to represent $\lambda_{H}$ to make our discussion more coherent. For this symmetric matrix of order four, we should convict eight different cases depending on the sign distributions of the off-diagonal elements to determine whether this matrix is copositive. In these cases, all the diagonal elements should be positive as a generic condition, and the eight cases are given as follows Chakrabortty:2013mha :
  • Case I: The matrix $\mathcal{S}$ is copositive if and only if $\lambda_{ii}\geq 0$ when all the off-diagonal elements are positive.
  • Case II: The matrix $\mathcal{S}$ is copositive if and only if $(\lambda_{ii}\lambda_{jj}-\lambda_{ij}^{2})\geq 0$ when $\lambda_{ij}\leq 0$ and other off-diagonal elements are positive.
  • Case III: The matrix $\mathcal{S}$ is copositive if and only if $(\lambda_{ii}\lambda_{jj}-\lambda_{ij}^{2})\geq 0,(\lambda_{ll}\lambda_{kk}-\lambda_{lk}^{2})\geq 0$ when $\lambda_{ij},\lambda_{lk}\leq 0$ and other off-diagonal elements are positive.
  • Case IV: When $\lambda_{ij},\lambda_{ik}\leq 0$, we should have $(\lambda_{ii}\lambda_{jk}-\lambda_{ij}\lambda_{ik}+\sqrt{(\lambda_{ii}\lambda_{jj}-\lambda_{ij}^{2})(\lambda_{ii}\lambda_{kk}-\lambda_{ik}^{2})})\\ \geq 0$ to make sure the matrix $\mathcal{S}$ is copositive.
  • Case V: The matrix $\mathcal{S}$ is copositive if and only if the following order three matrix is copositive:
    $\left(\begin{array}[]{ccc}\lambda_{ii}&\lambda_{ij}&\lambda_{ik}\\ &\lambda_{jj}&\lambda_{jk}\\ &&\lambda_{kk}\\ \end{array}\right).$
    if $\lambda_{ij},\lambda_{jk},\lambda_{ik}\leq 0$ while the other off-diagonal elements are positive.
  • Case VI: The matrix $\mathcal{S}$ is copositive if the following matrix is copositive:
    $\left(\begin{array}[]{ccc}\lambda_{ii}\lambda_{jj}-\lambda_{ij}^{2}&\lambda_{ii}\lambda_{jk}-\lambda_{ij}\lambda_{ik}&\lambda_{ii}\lambda_{jl}-\lambda_{ij}\lambda_{il}\\ &\lambda_{ii}\lambda_{kk}-\lambda_{ik}^{2}&\lambda_{ii}\lambda_{kl}-\lambda_{ik}\lambda_{il}\\ &&\lambda_{ii}\lambda_{ll}-\lambda_{il}^{2}\\ \end{array}\right).$
    when $\lambda_{ij},\lambda_{ik},\lambda_{il}\leq 0$ and other off-diagonal elements are positive.
  • Case VII: The following matrix of order three should be copositive to make sure $\mathcal{S}$ is copositive:
    $\left(\begin{array}[]{ccc}\lambda_{kk}\big{(}\lambda_{jj}\lambda_{ik}^{2}-2\lambda_{ij}\lambda_{ik}\lambda_{jk}+\lambda_{ii}\lambda_{jk}^{2}\big{)}&\lambda_{kk}(\lambda_{jj}\lambda_{ik}-\lambda_{ij}\lambda_{jk})&\lambda_{kk}(\lambda_{ik}\lambda_{jl}-\lambda_{jk}\lambda_{il})\\ &\lambda_{jj}\lambda_{kk}-\lambda_{jk}^{2}&\lambda_{kk}\lambda_{jl}-\lambda_{jk}\lambda_{kl}\\ &&\lambda_{kk}\lambda_{ll}-\lambda_{kl}^{2}\\ \end{array}\right).$
    when $\lambda_{ij},\lambda_{jk},\lambda_{kl}\leq 0$ and other off-diagonal elements are positive.
  • Case VIII: The following matrix of order three should be copositive to make sure $\mathcal{S}$ is copositive:
    $\left(\begin{array}[]{ccc}\lambda_{ll}(\lambda_{ii}\lambda_{jl}^{2}-2\lambda_{ij}\lambda_{il}\lambda_{jl}+\lambda_{jj}\lambda_{il}^{2})&\lambda_{ll}(\lambda_{ii}\lambda_{jl}-\lambda_{ij}\lambda_{il})&\lambda_{ll}(\lambda_{ik}\lambda_{jl}-\lambda_{il}\lambda_{jk})\\ &\lambda_{ii}\lambda_{ll}-\lambda_{il}^{2}&\lambda_{ll}\lambda_{ik}-\lambda_{il}\lambda_{kl}\\ &&\lambda_{kk}\lambda_{ll}-\lambda_{kl}^{2}\\ \end{array}\right).$
    when $\lambda_{ij},\lambda_{jk},\lambda_{kl},\lambda_{il}\leq 0$ and other off-diagonal elements are positive.
We should discuss these eight cases separately and thus will have different allowed parameter space from the view of theoretical constraint.
4 Dark matter phenomenology
In our model, there are two scalar dark matter particles stabilized by $Z_{2}\times Z_{2}^{\prime}$, which are related to SM particles with Higgs-portal interactions. The current observed dark matter relic density given by the Planck collaboration is $\Omega_{DM}h^{2}=0.1198\pm 0.0012$ Belanger:2020gnr , and we consider dark matter production in our model to be generated with the “Freeze-out” mechanism. Both $S_{1}$ and $S_{2}$ will contribute to dark matter relic density and the Boltzmann equations for the number density of $S_{1}$ and $S_{2}$ are given as follows:
$\frac{dn_{1}}{dt}+3Hn_{1}$ $=$ $-\langle\sigma v\rangle^{S_{1}S_{1}^{*}\to XX}(n_{1}^{2}-{\bar{n_{1}}}^{2})-\langle\sigma v\rangle^{S_{1}S_{1}^{*}\to h_{1,2}h_{1,2}}(n_{1}^{2}-{\bar{n_{1}}}^{2})$
-- $\langle\sigma v\rangle^{S_{1}S_{1}^{*}\to S_{2}S_{2}^{*}}(n_{1}^{2}-n_{2}^{2}\frac{\bar{n_{1}}^{2}}{\bar{n_{2}}^{2}})$
$\frac{dn_{2}}{dt}+3Hn_{2}$ $=$ $-\langle\sigma v\rangle^{S_{2}S_{2}^{*}\to XX}(n_{2}^{2}-{\bar{n_{2}}}^{2})-\langle\sigma v\rangle^{S_{2}S_{2}^{*}\to h_{1,2}h_{1,2}}(n_{2}^{2}-{\bar{n_{2}}}^{2})$ (16)
-- $\langle\sigma v\rangle^{S_{2}S_{2}^{*}\to S_{1}S_{1}^{*}}(n_{2}^{2}-n_{1}^{2}\frac{\bar{n_{2}}^{2}}{\bar{n_{1}}^{2}})$
where $n_{1}(n_{2})$ is the number density of $S_{1}(S_{2})$, $\bar{n_{1}}(\bar{n_{2}})$ is the number density in thermal equilibrium, $H$ is the Hubble expansion rate of the Universe, $X$ denotes SM particles and $\langle\sigma v\rangle$ is the thermally averaged annihilation cross section. To calculate the DM relic density numerically we use the micrOMGEAs 5.0.6 package Alguero:2022inz , in which the model has been implemented through the FeynRules package Alloul:2013bka .
On the other hand, dark matter can scatter off the nuclei with the Higgs-mediated t-channel processes, which should be constrained by direct detection results. The effective Lagrangian related to DM-quark elastic scattering can be given by Basak:2021tnj :
$\mathcal{L}_{q,eff}=\sum_{S=S_{1},S_{2}}-\frac{m_{q}}{2v_{0}}(\frac{\mathcal{C}_{h_{1}SS}}{m_{h_{1}}^{2}}+\frac{\mathcal{C}_{h_{2}SS}}{m_{h_{2}}^{2}})SS\bar{q}q$ (17)
where $\mathcal{C}_{h_{1}SS}$ as well as $\mathcal{C}_{h_{2}SS}$ are the couplings of dark matter $S_{1,2}$ with the $h_{1,2}$ and
$\mathcal{C}_{h_{1}S_{1}S_{1}}=-2i\cos\theta\lambda_{14}v_{0}+2i\sin\theta\lambda_{13}v_{1},$
$\mathcal{C}_{h_{2}S_{1}S_{1}}=-2i\sin\theta\lambda_{14}v_{0}-2i\cos\theta\lambda_{13}v_{1},$
$\mathcal{C}_{h_{1}S_{2}S_{2}}=-2i\cos\theta\lambda_{24}v_{0}+2i\cos\theta\lambda_{23}v_{1},$
$\mathcal{C}_{h_{2}S_{2}S_{2}}=-2i\sin\theta\lambda_{24}v_{0}-2i\cos\theta\lambda_{23}v_{1}$ (18)
Furthermore, the effective Lagrangian related to DM-nucleon elastic scattering can be given by:
$\mathcal{L}_{N,eff}=\sum_{S=S_{1},S_{2}}=\frac{m_{N}-\frac{7}{9}m_{B}}{v_{0}}(\frac{\mathcal{C}_{h_{1}SS}}{m_{h_{1}}^{2}}+\frac{\mathcal{C}_{h_{2}SS}}{m_{h_{2}}^{2}})SS\bar{N}N$ (19)
where $m_{N}$ represents the nucleon mass and $m_{B}$ is the baryon mass in the chiral limit He:2008qm . The expression of the cross-section for the spin-independent DM- nucleon elastic scattering can thus be given as follows:
$\sigma_{S_{i}N}=\frac{m_{N}^{4}f_{N}^{2}}{4\pi(m_{N}+m_{i})^{2}}\times(\frac{\mathcal{C}_{h_{1}S_{i}S_{i}}}{m_{h_{1}}^{2}}+\frac{\mathcal{C}_{h_{2}S_{i}S_{i}}}{m_{h_{2}}^{2}})^{2},(i=1,2)$ (20)
where $f_{N}$ is the Higgs-nucleon Form factor with $f_{N}=0.308(18)$ according to the phenomenological and lattice-QCD calculations Bell:2013wua . Since we have two dark matter particles in the model, we should compare the value of $\xi_{i}\sigma_{S_{i}N}$ against the direct detection constraints provided by the experiment results instead of the cross-section, where $\xi_{i}$ is the fraction of $S_{i}$ in the total relic density defined by:
$\xi_{i}=\frac{\Omega_{i}}{\Omega_{DM}},(i=1,2)$ (21)
and $\Omega_{DM}$ is the observed value of the dark matter density $\Omega_{i}$ is the density of $S_{i}(i=1,2)$. In our analysis, we will consider the current direct detection limit set by the XENONnT XENON:2015gkh and LZ LZ:2022lsv .
5 Discussion
In this part, we estimate the parameter space from the view of copositive criteria and dark matter phenomenology related to relic density and direct detection constraints. As we mentioned above, there are eight cases related to copositive criteria that depend on the sign of the parameters, and each case corresponds to a viable parameter space that should be limited by the relic density constraint and direct detection constraint. On the other hand, regardless of the symmetric matrix $\mathcal{S}$ with the permutation $\lambda_{ij}=\lambda_{ji}$, we can still have different results when the indexes $i,j,k,l$ takes different permutations even just for one case. Hence, discussion about the parameter space will be complex for different permutations of the indexes and one should consider all the possible cases to obtain the viable parameter space.
5.1 Possible permutations of the indexes
To decrease the complexity and simplify the discussion, we consider the following assumptions. Firstly, notice that we have an exchange symmetry of $\lambda_{13}\leftrightarrow\lambda_{23}$, $\lambda_{14}\leftrightarrow\lambda_{24}$ and $S_{1}\leftrightarrow S_{2}$ due to the $Z_{2}\times Z_{2}^{\prime}$ symmetry, and there is no preference for the permutations of the indexes when considering the copositive criteria. Therefore, we will only consider one possible case when discussing the related parameters since we will have a similar shape of the parameter space. On the other hand, $\lambda_{34}$ is the coupling of the new Higgs particles, which makes little difference in dark matter results, and we omit it as a free parameter and will not consider the cases including the sign of $\lambda_{34}$ in the following discussion.
Taking the above assumptions into consideration, we can conclude all the possible cases as follows when discussing copositive criteria:
I) For Case I with $\lambda_{ij}\geqslant 0$, we have all the couplings larger than 0.
II) For case II with $\lambda_{ij}\leqslant 0$, we have the following permutations: $\lambda_{12}\leqslant 0$ or $\lambda_{13}\leqslant 0$ or $\lambda_{24}\leqslant 0$ or $\lambda_{14}\leqslant 0$ or $\lambda_{23}\leqslant 0$. As we mentioned above, we ignore the case $\lambda_{34}\leqslant 0$. Meanwhile, for the cases $\lambda_{13}\leqslant 0$ and $\lambda_{23}\leqslant 0$, we only need to consider one of these two cases due to the exchange symmetry and here we can consider $\lambda_{13}\leqslant 0$ for the following discussion. Similarly, for $\lambda_{14}\leqslant 0$ and $\lambda_{24}\leqslant 0$ we consider $\lambda_{24}\leqslant 0$. Hence, there are 3 cases to be discussed for Case II with $\lambda_{12}\leqslant 0$ or $\lambda_{13}\leqslant 0$ or $\lambda_{14}\leqslant 0$.
III) For Case III with $\lambda_{ij},\lambda_{lk}\leqslant 0$, there are $C^{2}_{4}=3$ kinds of permutations in total, but we will only need to discuss the case with $\lambda_{14},\lambda_{23}\leqslant 0$ and the case $\lambda_{12},\lambda_{34}\leqslant 0$ is ignored while $\lambda_{13},\lambda_{24}\leqslant 0$ has the similar results with $\lambda_{14},\lambda_{23}\leqslant 0$.
IV) For Case IV with $\lambda_{ij},\lambda_{ik}\leqslant 0$, the amount of the possible permutations is $\frac{4\times 3\times 2}{2}-2\times 2=8$ without the cases including the sign of $\lambda_{34}$, and we will have 5 different kinds of permutations in total after taking exchange symmetry into consideration. Concretely speaking, we will discuss the cases with $\lambda_{12},\lambda_{13}\leqslant 0$,$\lambda_{12},\lambda_{14}\leqslant 0$,$\lambda_{13},\lambda_{14}\leqslant 0$,$\lambda_{23},\lambda_{13}\leqslant 0$,$\lambda_{14},\lambda_{24}\leqslant 0$
V) For Case V with $\lambda_{ij},\lambda_{jk},\lambda_{ik}\leqslant 0$, we have $\frac{4\times 3\times 2}{2\times 2\times 2}=3$ kinds of different permutations, and we have $\lambda_{12},\lambda_{23},\lambda_{13}\leqslant 0$ as well as $\lambda_{12},\lambda_{24},\lambda_{14}\leqslant 0$ after ignoring the case including $\lambda_{34}$.
VI) For Case VI with $\lambda_{ij},\lambda_{ik},\lambda_{il}\leqslant 0$ we have $\lambda_{12},\lambda_{13},\lambda_{14}\leqslant 0$ after taking exchange symmetry into consideration.
VII) For Case VII with $\lambda_{ij},\lambda_{jk},\lambda_{kl}\leqslant 0$, we have 3 different kinds of permutations left after taking exchange symmetry into consideration and ignoring the case including $\lambda_{34}$. Concretely speaking, we have $\lambda_{13},\lambda_{23},\lambda_{24}\leqslant 0$, $\lambda_{14},\lambda_{24},\lambda_{23}\leqslant 0$ and $\lambda_{12},\lambda_{13},\lambda_{24}\leqslant 0$
VIII) For Case VIII with $\lambda_{ij},\lambda_{jk},\lambda_{kl},\lambda_{il}\leqslant 0$, we have $\frac{4\times 3\times 2}{2\times 4}=3$ kinds of permutations, and we only need to discuss the case $\lambda_{13},\lambda_{23},\lambda_{24},\lambda_{14}\leqslant 0$ after considering exchange symmetry and ignoring the case including $\lambda_{34}$.
In conclusion, we will consider $17(1+3+1+5+2+1+3+1)$ different kinds of permutations in total in the following discussion.
5.2 Theortical constraint of copositive criteria
Following the above discussion, we consider the 17 different cases with theoretical constraint of copositive criteria on the viable parameter space. In this work, we focus on the parameters related to dark matter phenomenology and copositive criteria. We can choose the following parameters as inputs:
$v_{1},\sin\theta,m_{h_{2}},m_{1,2},\lambda_{11},\lambda_{12},\lambda_{13},\lambda_{14},\lambda_{22},\lambda_{23},\lambda_{24}$ (22)
while $\lambda_{33}$,$\lambda_{34}$ and $\lambda_{44}(\lambda_{H})$ can be given by Eq. 2. Note that for the large couplings, the parameter space can be excluded by direct detection constraints, and in the following discussion we assume the couplings to be smaller than $\pi$ for simplicity. On the other hand, since the couplings $\lambda_{ii}$ are less relevant with dark matter phenomenology with $i=1,2,3,4$ and one can always fine-tune these couplings to satisfy the related inequalities to guarantee the copositive criteria.
For Case I, Case II and Case III, the copositive criteria can be satisfied by fine-tuning the $\lambda_{ii}$ with $i=1,2,3,4$ as we mentioned above. For Case IV, the parameter space is constrained by the inequality to satisfy the copositive criteria. In Fig. 1(a), we give the result of Case IV, where the colored region is the viable parameter space of $\lambda_{ij}-\lambda_{ik}$ satisfying the copositive criteria and we have set $\lambda_{ii}=\lambda_{jj}=\lambda_{kk}=\pi$ and $\lambda_{jk}=1$. The small negative $\lambda_{ij}-\lambda_{ik}$ pair is excluded for the copositive criteria and for a smaller $\lambda_{jk}$ the allowed parameter space for $(\lambda_{ij},\lambda_{ik})$ is more stringent according to copositive criteria.
According to Case V, the copositive criteria demands the symmetric matrix of order three is copositive, and we have:
$\lambda_{ij}+\sqrt{\lambda_{ii}\lambda_{jj}}\geqslant 0,\lambda_{ik}+\sqrt{\lambda_{ii}\lambda_{kk}}\geqslant 0,\lambda_{jk}+\sqrt{\lambda_{jj}\lambda_{kk}}\geqslant 0,\sqrt{\lambda_{ii}\lambda_{jj}\lambda_{kk}}+\lambda_{ij}\sqrt{\lambda_{kk}}+\lambda_{ik}\sqrt{\lambda_{jj}}$
$+\lambda_{jk}\sqrt{\lambda_{ii}}+\sqrt{2(\lambda_{ij}+\sqrt{\lambda_{ii}\lambda_{jj}})(\lambda_{ik}+\sqrt{\lambda_{ii}\lambda_{kk}})(\lambda_{jk}+\sqrt{\lambda_{jj}\lambda_{kk}})}\geqslant 0$ (23)
Figure 1: The allowed parameter space constrained by the copositive criteria corresponds to the colored region, where (a) represents the result of Case IV, (b) corresponds to the result of Case V, (c) corresponds to Case VI, (d) is the result of Case VII and (e) corresponds to Case VIII.
We give the result of Case V in Fig. 1(b), and the colored region is the viable parameter space of $\lambda_{ij}-\lambda_{ik}$ satisfying the copositive criteria, where we have set $\lambda_{ii}=\lambda_{jj}=\lambda_{kk}=\pi$ and $\lambda_{jk}=-1$. For the negative $\lambda_{ij}-\lambda_{ik}$, the left-bottom region corresponding to small $\lambda_{ij}-\lambda_{ik}$ is excluded with the copositive criteria.
For Case VI, the copositive criteria demands that:
$\lambda_{ii}\lambda_{jk}-\lambda_{ij}\lambda_{ik}+\sqrt{(\lambda_{ii}\lambda_{jj}-\lambda_{ij}^{2})(\lambda_{ii}\lambda_{kk}-\lambda_{ik}^{2})}\geqslant 0,$ (24)
$\lambda_{ii}\lambda_{jl}-\lambda_{ij}\lambda_{il}+\sqrt{(\lambda_{ii}\lambda_{jj}-\lambda_{ij}^{2})(\lambda_{ii}\lambda_{ll}-\lambda_{il}^{2})}\geqslant 0,$ (25)
$\lambda_{ii}\lambda_{kl}-\lambda_{ik}\lambda_{il}+\sqrt{(\lambda_{ii}\lambda_{kk}-\lambda_{ik}^{2})(\lambda_{ii}\lambda_{ll}-\lambda_{il}^{2})}\geqslant 0,$ (26)
$\sqrt{(\lambda_{ii}\lambda_{jj}-\lambda_{ij}^{2})(\lambda_{ii}\lambda_{kk}-\lambda_{ik}^{2})(\lambda_{ii}\lambda_{ll}-\lambda_{il}^{2})}+(\lambda_{ii}\lambda_{jk}-\lambda_{ij}\lambda_{ik})\sqrt{\lambda_{ii}\lambda_{ll}-\lambda_{il}^{2}}$
$+(\lambda_{ii}\lambda_{jl}-\lambda_{ij}\lambda_{il})\sqrt{\lambda_{ii}\lambda_{kk}-\lambda_{ik}^{2}}+(\lambda_{ii}\lambda_{kl}-\lambda_{ik}\lambda_{il})\sqrt{\lambda_{ii}\lambda_{jj}-\lambda_{ij}^{2}}+\sqrt{2}$
$\sqrt{(\lambda_{ii}\lambda_{jl}-\lambda_{ij}\lambda_{il})+\sqrt{(\lambda_{ii}\lambda_{jj}-\lambda_{ij}^{2})(\lambda_{ii}\lambda_{kk}-\lambda_{ik}^{2})}}\sqrt{(\lambda_{ii}\lambda_{jl}-\lambda_{ij}\lambda_{il}+\sqrt{(\lambda_{ii}\lambda_{jj}-\lambda_{ij}^{2})(\lambda_{ii}\lambda_{ll}-\lambda_{il}^{2})}}$
$\sqrt{(\lambda_{ii}\lambda_{kl}-\lambda_{ik}\lambda_{il})+\sqrt{(\lambda_{ii}\lambda_{kk}-\lambda_{ik}^{2})(\lambda_{ii}\lambda_{ll}-\lambda_{il}^{2})}}\geqslant 0$ (27)
The parameter space is constrained by the above inequalities, and we give the result of $\lambda_{ij}-\lambda_{ik}$ satisfying copositive criteria in Fig. 1(c), where we set $\lambda_{ii}=\lambda_{jj}=\lambda_{kk}=\lambda_{ll}=\pi$,$\lambda_{jk}=\lambda_{jl}=0.1$, $\lambda_{kl}=1$ and $\lambda_{il}=-1$. The red and magenta region corresponds to the results of Eq. 27 and Eq. 24, while the green line represents the lower bound of Eq. 25 with the left region of the line being excluded. What’s more, one can estimate that Eq. 26 always holding for $-\pi\leqslant\lambda_{ij}\leqslant 0$. The intersection part of the colored region lying on the right side of the green line is the viable parameter space satisfying copositive criteria.
For Case VII, one can define:
$A_{11}=\lambda_{kk}(\lambda_{jj}\lambda_{ik}^{2}-2\lambda_{ij}\lambda_{ik}\lambda_{jk}+\lambda_{ii}\lambda_{jk}^{2})$
$A_{12}=\lambda_{kk}(\lambda_{jj}\lambda_{ik}-\lambda_{ij}\lambda_{jk}),~{}~{}A_{13}=\lambda_{kk}(\lambda_{ik}\lambda_{jl}-\lambda_{jk}\lambda_{il})$
$A_{22}=\lambda_{jj}\lambda_{kk}-\lambda_{jk}^{2},~{}~{}A_{23}=\lambda_{kk}\lambda_{jl}-\lambda_{jk}\lambda_{kl},~{}~{}A_{33}=\lambda_{kk}\lambda_{ll}-\lambda_{kl}^{2}$
the copositive criteria demands the symmetric matrix of order three is copositive, and we have:
$A_{11}\geqslant 0,A_{22}\geqslant 0,A_{33}\geqslant 0$ (28)
$A_{12}+\sqrt{A_{11}A_{22}}\geqslant 0,$ (29)
$A_{13}+\sqrt{A_{11}A_{33}}\geqslant 0,$ (30)
$A_{23}+\sqrt{A_{22}A_{33}}\geqslant 0$ (31)
$A_{13}\sqrt{A_{22}}+A_{23}\sqrt{A_{11}}+\sqrt{2(A_{12}+\sqrt{A_{11}A_{22}})(A_{13}+\sqrt{A_{11}A_{33}})(A_{23}+\sqrt{A_{22}A_{33}})}$
$+\sqrt{A_{11}A_{22}A_{33}}+A_{12}\sqrt{A_{33}}\geqslant 0,$ (32)
One can estimate that $\pi\lambda_{jl}^{2}-2\lambda_{ij}\lambda_{il}\lambda_{jl}+\pi\lambda_{il}^{2}\geqslant-\lambda_{ij}\lambda_{jl}^{2}-2\lambda_{ij}\lambda_{il}\lambda_{jl}-\lambda_{ij}\lambda_{il}^{2}=-\lambda_{ij}(\lambda_{jl}-\lambda_{il})^{2}\geqslant 0$ so that $A_{11}\geqslant 0$ always holds and we have $A_{22}\geqslant 0$ and $A_{33}\geqslant 0$ as we mentioned above. For Case VII, we give the result of $\lambda_{ij}-\lambda_{jk}$ in Fig. 1(d) where we set $\lambda_{ik}=\lambda_{jl}=0.1$,$\lambda_{il}=1$ and $\lambda_{kl}=-1$. The red,blue,green and magenta region correspond to the results of Eq. 29,Eq. 30,Eq. 31 and Eq. 32 respectively. The intersection part of these regions are the allowed parameter space that satisfying copositive criteria.
For Case VIII, one can define:
$B_{11}=\lambda_{ll}(\lambda_{ii}\lambda_{jl}^{2}-2\lambda_{ij}\lambda_{il}\lambda_{jl}+\lambda_{jj}\lambda_{il}^{2})$
$B_{12}=\lambda_{ll}(\lambda_{ii}\lambda_{jl}-\lambda_{ij}\lambda_{il}),B_{13}=\lambda_{ll}(\lambda_{ik}\lambda_{jl}-\lambda_{il}\lambda_{jk})$
$B_{22}=\lambda_{ii}\lambda_{ll}-\lambda_{il}^{2},B_{23}=\lambda_{ll}\lambda_{ik}-\lambda_{il}\lambda_{kl},B_{33}=\lambda_{kk}\lambda_{ll}-\lambda_{kl}^{2}$
Similarly, the copositive criteria demands that:
$B_{11}\geqslant 0,B_{22}\geqslant 0,B_{33}\geqslant 0$ (33)
$B_{12}+\sqrt{B_{11}B_{22}}\geqslant 0,$ (34)
$B_{13}+\sqrt{B_{11}B_{33}}\geqslant 0,$ (35)
$B_{23}+\sqrt{B_{22}B_{33}}\geqslant 0,$ (36)
$B_{13}\sqrt{B_{22}}+B_{23}\sqrt{B_{11}}+\sqrt{2(B_{12}+\sqrt{B_{11}B_{22}})(B_{13}+\sqrt{B_{11}B_{33}})(B_{23}+\sqrt{B_{22}B_{33}})}$
$+\sqrt{B_{11}B_{22}B_{33}}+B_{12}\sqrt{B_{33}}\geqslant 0$ (37)
One can estimate that $B_{11}\geqslant 0$ similar with $A_{11}\geqslant 0$ and also $B_{22}\geqslant 0$ as well as $B_{33}\geqslant 0$. For Case VIII, we give the result of $\lambda_{ij}-\lambda_{il}$ in Fig. 1(e) where we set $\lambda_{jk}=\lambda_{jl}=-1$,$\lambda_{kl}=1$ and $\lambda_{ij}=0.1$. The red, blue, green and magenta regions correspond to the results of Eq. 34,Eq. 35,Eq. 36 and Eq. 37 respectively. The intersection part of these regions is the allowed parameter space that satisfies copositive criteria.
In this part, we discuss the theoretical constraint on the parameter space from Case IV to Case VIII. For different copositive criteria, we will have different viable parameter spaces according to Fig. 1. We give a general discussion without considering index permutations index. However, when discussing dark matter phenomenology, different couplings will play different roles in determining relic density and be constrained by direct detection constraint, and we should consider the 17 different cases in total from Case I to Case VIII as we mentioned above.
5.3 Relic density and direct detection constraints
Following the above discussion, in this part, we consider relic density and direct detection constraints on the parameter space. As we have mentioned above, there are 17 different cases to be discussed when we consider copositive criteria. For simplicity, we fix $\sin\theta=0.01$, $m_{h_{2}}=1.5$ TeV,$v_{1}=2$ TeV and we randomly scan the parameter space with :
$m_{1,2}\subset[1\mathrm{TeV},4\mathrm{TeV}],|\lambda_{12}|\subset[0.0001,3.14],|\lambda_{13}|\subset[0.0001,3.14],$
$|\lambda_{14}|\subset[0.0001,3.14],|\lambda_{23}|\subset[0.0001,3.14],|\lambda_{24}|\subset[0.0001,3.14]$
We consider the copositive criteria, dark matter relic density constraint and direct detection constraint on the model and give the results of the 17 different cases from Fig. 2 to Fig. 10.
In Fig. 2, we give the results of Case I with $\lambda_{ij}\geqslant 0$, where Fig. 2(a) and Fig. 2(b) correspond to the viable parameter space $\lambda_{13}-\lambda_{23}$ as well as $\lambda_{14}-\lambda_{24}$ respectively and points with different colors represent $\lambda_{12}$ taking different values. According to Fig. 2(a), the allowed parameter space for $\lambda_{13}$ and $\lambda_{23}$ are constrained within $[0.1,3.14]$, while the smaller values are excluded for dark matter density being over-abundant. On the other hand, within the viable parameter space, a smaller $\lambda_{13}$ always demands a larger $\lambda_{23}$ to meet relic density constraint and vice due to the exchange symmetry of $S_{1}$ and $S_{2}$. With the increase of $\lambda_{13}$, the parameter space for $\lambda_{23}$ is more flexible. For $\lambda_{14}$ and $\lambda_{24}$, the allowed regions are $[10^{-3},3.14]$, which are less relevant according to Fig. 2(b). Similarly, the small $\lambda_{14}-\lambda_{24}$ region is excluded by relic density constraint, and a smaller $\lambda_{14}$ always demands a larger $\lambda_{24}$ to meet relic density constraint and verse vice within the viable parameter space. On the other hand, for the large $\lambda_{14}$ such as $\lambda_{14}>1$, the upper bound of $\lambda_{24}$ is constrained stringently with about $\lambda_{24}<0.7$ for dark matter density being under abundant and vice verse. For $\lambda_{12}$ , one can take value ranging from $[0.0001,3.14]$ according to Fig. 2.
Figure 2: Results of Case I with $\lambda_{ij}\geqslant 0$. The left picture corresponds to the viable parameter space of $\lambda_{13}-\lambda_{23}$ and the right picture is $\lambda_{14}-\lambda_{24}$, where points with different colors represent $\lambda_{12}$ taking different values in both pictures.
Figure 3: Results of Case II, where (a) and (d) are results of $\lambda_{12}\leqslant 0$, (b) and (e) are $\lambda_{13}\leqslant 0$, and (c) and (f) are $\lambda_{14}\leqslant 0$.
Figure 4: Results of Case III with $\lambda_{23}\leqslant 0,\lambda_{14}\leqslant 0$, where Fig. 4(a) is the viable parameter space of $(\lambda_{13}-|\lambda_{23}|)$ and Fig. 4(b) is the result of $(|\lambda_{14}|-\lambda_{24})$.
Figure 5: Viable parameter space of $(|\lambda_{13}|-|\lambda_{23}|)$ with Case IV, where (a),(b),(c),(d),(e) are results of $\lambda_{12},\lambda_{13}\leqslant 0$,$\lambda_{12},\lambda_{14}\leqslant 0$,$\lambda_{13},\lambda_{14}\leqslant 0$,$\lambda_{23},\lambda_{13}\leqslant 0$ and $\lambda_{14},\lambda_{24}\leqslant 0$ respectively.
Figure 6: Viable parameter space of $(|\lambda_{14}|-|\lambda_{24}|)$ with Case IV, where (a),(b),(c),(d),(e) are results of $\lambda_{12},\lambda_{13}\leqslant 0$,$\lambda_{12},\lambda_{14}\leqslant 0$,$\lambda_{13},\lambda_{14}\leqslant 0$,$\lambda_{23},\lambda_{13}\leqslant 0$ and $\lambda_{14},\lambda_{24}\leqslant 0$ respectively.
Figure 7: Results of Case V, where (a) and (c) are results of $\lambda_{12},\lambda_{23},\lambda_{13}\leqslant 0$, (b) and (d) are $\lambda_{12},\lambda_{24},\lambda_{14}\leqslant 0$.
Figure 8: Results of Case VI with $\lambda_{12},\lambda_{13},\lambda_{14}\leqslant 0$, where Fig. 8(a) is the viable parameter space of $(|\lambda_{13}|-|\lambda_{23}|)$ and Fig. 8(b) is the result of $(|\lambda_{14}|-|\lambda_{24}|)$.
We give the results of Case II in Fig. 3, where Fig. 3(a) and Fig. 3(d) are results of $\lambda_{12}\leqslant 0$, Fig. 3(b) and Fig. 3(e) are $\lambda_{13}\leqslant 0$, and Fig. 3(c) and Fig. 3(f) are $\lambda_{14}\leqslant 0$. According to Fig. 3(a),(b),(c), we give the viable parameter space of $(|\lambda_{13}|,|\lambda_{23}|)$ while in Fig. 3(d),(e),(f), we show the results of $(|\lambda_{14}|,|\lambda_{24}|)$. As we mentioned above, for Case II, the parameter space can just be constrained by relic density and direct detection constraints and rid of the copositive criteria. Since the cross section of $S_{i}S_{i}\to h_{i}h_{i}(i=1,2)$ is determined by Higgs-mediated processes, the quartic term $|S_{i}|^{2}|h_{i}|^{2}$ and the t-channel processes exchange by dark matter, where the sign of the couplings $\lambda_{i3}(\lambda_{i4})$ with $i=1,2$ can make difference in the interference terms of these diagrams. On the other hand, the combined contribution of the Higgs-mediated processes and the quartic term $|S_{1}|^{2}|S_{2}|^{2}$ will determine the cross section of $S_{1}S_{1}\to S_{2}S_{2}$, where the sign of $\lambda_{12}$ can also play an important role in the result. However, such interactions between $S_{1}$ and $S_{2}$ just adjust the relative relic density of $S_{1}$ and $S_{2}$, and we will not focus on the influence of $\lambda_{12}$ in this work due to the exchange symmetry. Note that we have chosen the small $\sin\theta$ so that the contribution of $S_{i}S_{i}\to h_{1}h_{2}(i=1,2)$ are suppressed. Due to the different choices for the signs of the couplings, the shapes of the viable parameter space for the three cases are slightly different according to Fig. 3. The expression of these processes in the limit of $\sin\theta\to 0$ can be found in App. A.4. Particularly, for $\lambda_{14}\leqslant 0$, the allowed region for $\lambda_{24}$ is about larger than $0.003$ due to the interference effect unlike another two cases, where $\lambda_{24}$ can take value ranging from $[10^{-3},3.14]$.
In Fig. 4, we give the results of Case III with $\lambda_{23}\leqslant 0,\lambda_{14}\leqslant 0$, where the left picture is the viable parameter space of $(\lambda_{13},|\lambda_{23}|)$ and the right picture is the result of $(|\lambda_{14}|,\lambda_{24})$. In this case, one can rid of the copositive criteria by fine-tuning so that the main limits on the parameter space arise from the relic density and direct detection constraints. Similar to the above cases, one of the features of the shape of the viable parameter space for $(\lambda_{13},|\lambda_{23}|)$ is that the points all lie in the right-top region while for $(|\lambda_{14}|,\lambda_{24})$ such region is excluded. On the other hand, due to the interference effect, the allowed parameter space for the couplings $\lambda_{13}$,$\lambda_{14}$, $\lambda_{23}$ and $\lambda_{24}$ are slightly different depending on the sign of these parameters.
In Fig. 5 and Fig. 6, we give the results of Case IV, where picture (a) to (e) in Fig. 5 correspond to the viable parameter spaces of $(|\lambda_{13}|-|\lambda_{23}|)$ with $\lambda_{12},\lambda_{13}\leqslant 0$, $\lambda_{12},\lambda_{14}\leqslant 0$,$\lambda_{13},\lambda_{14}\leqslant 0$,$\lambda_{23},\lambda_{13}\leqslant 0$ and $\lambda_{14},\lambda_{24}\leqslant 0$. Similarly, we give the results of $(|\lambda_{14}|,|\lambda_{24}|)$ with different cases respectively in Fig. 6. According to Fig. 5, the parameter space of $(|\lambda_{13}|,|\lambda_{23}|)$ is limited stringently by the relic density constraint and direct detection constraint like the above cases. There is a slight difference between $\lambda_{23},\lambda_{13}\leqslant 0$ and other cases, where for the larger $(|\lambda_{13}|,|\lambda_{14}|)$, the parameter space is excluded for dark matter density according to Fig. 5(d), which arises from the fact that dark matter density is under abundance due to the large interference effect. According to Fig. 6, the allowed parameter spaces for $(|\lambda_{14}|,|\lambda_{24}|)$ are also different for the five different cases. There is no preference for $\lambda_{14}$ and $\lambda_{24}$ because of the exchange symmetry, and the viable parameter space of $(|\lambda_{14}|,|\lambda_{24}|)$ are almost symmetric according to Fig. 6(a),Fig. 6(c),Fig. 6(d) and Fig. 6(e). However, when we consider the copositive criteria, the sign of the parameters should be discussed and the allowed $|\lambda_{14}|-|\lambda_{24}|$ values are slightly different due to the interference effect according to Fig. 6(b).
We give the results of Case V in Fig. 7, where pictures (a) and (c) correspond to the viable parameter spaces of $\lambda_{12},\lambda_{23},\lambda_{13}\leqslant 0$, while (b) and (d) are $\lambda_{12},\lambda_{24},\lambda_{14}\leqslant 0$. Unlike the above cases where most of the points lying within the left-top region of $(|\lambda_{13}|,|\lambda_{23}|)$, the larger $|\lambda_{13}|$ and $|\lambda_{23}|$ are excluded by the relic density constraint when we consider the signs of the couplings according to Fig. 7(a). For $\lambda_{14},\lambda_{24},\lambda_{12}\leqslant 0$, the interference effect can make difference on the parameter space of $(|\lambda_{14}|,|\lambda_{24}|)$ according to Fig. 7(d), where the shape for the allowed value $|\lambda_{14}|-|\lambda_{24}|$ is slightly not symmetric. On the other hand, the viable value for $|\lambda_{13}|-|\lambda_{23}|$ and $|\lambda_{14}|-|\lambda_{24}|$ are similar with the above cases according to Fig. 7(b) and Fig. 7(d) when the choices of the signs of the couplings make little difference on the parameter space.
Figure 9: Results of Case VII, where (a) and (d) are the results of $\lambda_{13},\lambda_{23},\lambda_{24}\leqslant 0$, (b) and (e) are results of $\lambda_{14},\lambda_{24},\lambda_{12}\leqslant 0$, and (c) and (f) are results of $\lambda_{12},\lambda_{13},\lambda_{24}\leqslant 0$.
Figure 10: Results of Case VIII with $\lambda_{13},\lambda_{23},\lambda_{24},\lambda_{14}\leqslant 0$.
We give the results of Case VI in Fig. 8 with $\lambda_{12},\lambda_{13},\lambda_{14}\leqslant 0$, where the left picture corresponds to the viable parameter space of $|\lambda_{13}|-|\lambda_{23}|$ and the right picture is $|\lambda_{14}|-|\lambda_{24}|$. The main feature of this case is that the upper bound for $(|\lambda_{14}|,|\lambda_{24}|)$ increases to about 0.6 due to the interference effect when considering the signs of the couplings.
We give the results of Case VII in Fig. 9 and Case VIII in Fig. 10, where Fig. 9(a) and Fig. 9(d) are the results of $\lambda_{13},\lambda_{23},\lambda_{24}\leqslant 0$, Fig. 9(b) and Fig. 9(e) are results of $\lambda_{14},\lambda_{24},\lambda_{12}\leqslant 0$, and Fig. 9(c) and Fig. 9(f) are results of $\lambda_{12},\lambda_{13},\lambda_{24}\leqslant 0$. According to Fig. 9(a), the larger $(|\lambda_{13}|,|\lambda_{23}|)$ are excluded by dark matter relic density constraint for $\lambda_{13},\lambda_{23},\lambda_{24}\leqslant 0$ while for the other two cases, the larger values of $|\lambda_{13}|$ and $|\lambda_{23}|$ can rid from the constraints due to the interference effect. Similar to the above-mentioned cases, the shapes of the parameter space of $(|\lambda_{14}|,|\lambda_{24}|)$ are slightly different for the three cases according to Fig. 9(d) to Fig. 9(f). For Case VIII, we have $\lambda_{13},\lambda_{23},\lambda_{24},\lambda_{14}\leqslant 0$, and Fig. 10(a) shows the viable parameter space of $(|\lambda_{13}|,|\lambda_{23}|)$ while Fig. 10(b) is the result of $(|\lambda_{14}|,|\lambda_{24}|)$.
In conclusion, according to Fig. 2 to Fig. 10, when we consider the signs of the couplings, we will have different viable parameter spaces under the different copositive criteria. We consider the 17 different cases and randomly scan the chosen parameter space under dark matter relic density constraint and direct detection constraint. For $(|\lambda_{13}|,|\lambda_{23}|)$, most of the points lie in the right-top region ranging from about 0.1 to 3.14 for most cases such as Case I , Case II and Case VI , where we have similar shapes of the parameter space no matter which couplings are negative. For other Cases, the choice of different couplings to be negative will make a difference in the parameter space. Particularly, for $\lambda_{13},\lambda_{23}\leqslant 0$ of Case IV, $\lambda_{12},\lambda_{13},\lambda_{23}\leqslant 0$ of Case V, $\lambda_{13},\lambda_{23},\lambda_{24}\leqslant 0$ of Case VII, $\lambda_{13},\lambda_{23},\lambda_{14},\lambda_{24}\leqslant 0$ of Case VIII, the larger $|\lambda_{13}|$ and $|\lambda_{23}|$ are excluded. In other words, the parameter space is more constrained as long as $\lambda_{13}\leqslant 0$ and $\lambda_{23}\leqslant 0$ for the copositive criteria. The reason is that the processes $S_{i}S_{i}\to h_{i}h_{i}(i=1,2)$ are efficiently influenced by the interference effect in the case of $\lambda_{13}\leqslant 0$ and $\lambda_{23}\leqslant 0$, and dark matter relic density will be under-abundant due to the larger annihilation processes arising from the larger couplings. For $(|\lambda_{14}|,|\lambda_{24}|)$, the interference effect on the cross section is less efficient compared with the for the low Higgs mass and SM vev, and the shapes of the parameter space of $(|\lambda_{14}|,|\lambda_{24}|)$ are almost similar among the different cases according to Fig. 2 to Fig. 10. We focus on the copositive criteria on the quartic couplings and do not show the viable parameter space of the masses of the scalars since for the TeV scale dark matter mass, the viable parameter space satisfying dark matter constrained is more flexible.
6 Summary and Outlook
Weakly interacting massive particles as dark matter are facing serious challenges that no signals have been found in the direct detection experiments so far, and one solution to such a problem is multi-component dark matter models, which include two or more dark matter species. On the other hand, the multi-component dark matter should be constrained by theoretical constraints and experiment constraints. Particularly, for the extended scalar dark matter models. the vacuum stability demand the scalar potential has to be bounded from below, which can put stringent constraints on the parameter space. What’s more, the copositive criteria allow to the derivation analytic necessary and sufficient vacuum stability conditions for the couplings, and different copositivity criteria will contribute to different viable parameter spaces depending on the choice of the signs of the couplings.
We consider a two-component scalar dark matter where dark matter particles are stabilized by extra $Z_{2}\times Z_{2}^{\prime}$ in this work. To obtain a stable vacuum, the $4\times 4$ matrix of the quadratic couplings should satisfy copositivity criteria. We pick up 17 different cases and focus on the quartic couplings in the model from the point of copositive criteria. The signs of the parameters will not only demand different copositivity criteria but can also influence the results of the cross section of the $2\to 2$ processes so that the dark matter relic density is due to the interference effect. We randomly scan the chosen parameter space with the copositivity criteria, dark matter relic density constraint and direct detection constraint. Among the 17 different cases, the shapes of the viable parameter space $|\lambda_{14}-|\lambda_{24}|$ are almost similar while for $|\lambda_{13}|-|\lambda_{23}|$ we can have different results due to the large interference effect of the $2\to 2$ processes of dark matter annihilation into scalars. As long as $\lambda_{13}\leqslant 0$ and $\lambda_{23}\leqslant 0$, the large value of $|\lambda_{13}|$ and $|\lambda_{23}|$ are excluded for the too large cross section so that dark matter density being under-abundant.
As we can see in this work, the different choices of the signs of couplings can indeed contribute to different parameter spaces with copositivity criteria. Our model is though simple while the complete discussion about the copositive criteria is complex for the different permutations of the indexes. For a more complex model, copositivity allows us to find analytic necessary and sufficient vacuum stability conditions so that estimate the parameter space theoretically.
Acknowledgements.
Hao Sun is supported by the National Natural Science Foundation of China (Grant No. 12075043, No. 12147205).
Appendix A Appendix
A.1 Perturbativity
To ensure the perturbative model, the contribution from loop correction should be smaller than the tree level values, and such constraints can be ensured with
$|\lambda_{13}|<4\pi,|\lambda_{23}|<4\pi,|\lambda_{12}|<4\pi,|\lambda_{14}|<4\pi,|\lambda_{24}|<4\pi.$ (38)
A.2 Perturbativity unitarity
The unitarity conditions come from the tree-level scalar-scalar scattering matrix which is dominated by the quartic contact interaction. The s-wave scattering amplitudes should lie under the perturbative unitarity limit, given the requirement the eigenvalues of the S-matrix $\mathcal{M}$ must be less than the unitarity bound given by $|\mathrm{Re}\mathcal{M}|<\frac{1}{2}$.
A.3 Renormalization Group Equations of the quartic couplings
In this part, we give the renormalization Group Equations (RGEs) of the model with SARAHStaub:2015kfa at one-loop level, and the beta functions of the quartic couplings are given by:
$\beta_{\lambda_{24}}$ $=+2\lambda_{12}\lambda_{14}-\frac{9}{10}g_{1}^{2}\lambda_{24}-\frac{9}{2}g_{2}^{2}\lambda_{24}+8\lambda_{22}\lambda_{24}+4\lambda_{24}^{2}+2\lambda_{23}\lambda_{34}+12\lambda_{24}\lambda_{44}+6\lambda_{24}y_{t}^{2}$ (39)
$\beta_{\lambda_{14}}$ $=-\frac{9}{10}g_{1}^{2}\lambda_{14}-\frac{9}{2}g_{2}^{2}\lambda_{14}+8\lambda_{11}\lambda_{14}+4\lambda_{14}^{2}+2\lambda_{12}\lambda_{24}+2\lambda_{13}\lambda_{34}+12\lambda_{14}\lambda_{44}+6\lambda_{14}y_{t}^{2}$ (40)
$\beta_{\lambda_{12}}$ $=2\lambda_{13}\lambda_{23}+4\lambda_{12}^{2}+4\lambda_{14}\lambda_{24}+8\lambda_{11}\lambda_{12}+8\lambda_{12}\lambda_{22}$ (41)
$\beta_{\lambda_{23}}$ $=2\lambda_{12}\lambda_{13}+4\Big{(}2\lambda_{22}\lambda_{23}+2\lambda_{23}\lambda_{33}+\lambda_{24}\lambda_{34}+\lambda_{23}^{2}\Big{)}$ (42)
$\beta_{\lambda_{13}}$ $=2\lambda_{12}\lambda_{23}+4\lambda_{13}^{2}+4\lambda_{14}\lambda_{34}+8\lambda_{11}\lambda_{13}+8\lambda_{13}\lambda_{33}$ (43)
where $y_{t}$ is the Top Yukawa coupling, $g_{1}$ and $g_{2}$ are the gauge couplings of $U(1)$ and $SU(2)$.
A.4 Cross section of $S_{i}S_{i}\to h_{i}h_{i}$ and $S_{i}S_{i}\to h_{1}h_{2}(i=1,2)$
We give the cross section of $S_{i}S_{i}\to h_{i}h_{i}$ and $S_{i}S_{i}\to h_{1}h_{2}$ $(i=1,2)$ in the part. We calculate the cross section with Calchep Belyaev:2012qa , and the expressions in the limit of $\sin\theta\to 0$ are given as follows.
$\sigma_{S_{1}S_{1}\to h_{1}h_{1}}$ $=\frac{\lambda_{14}^{2}}{8\pi s(s-4m_{1}^{2})}(\sqrt{(s-4m_{h_{1}}^{2})(s-4m_{1}^{2})}(\frac{8\lambda_{14}^{2}v_{0}^{4}}{m_{h_{1}}^{4}-4m_{h_{1}}^{2}m_{1}^{2}+m_{1}^{2}s}+\frac{(2m_{h_{1}}^{2}+s)^{2}}{(m_{h_{1}}^{2}-s)^{2}})$
$+$ $\frac{8\lambda_{14}v_{0}^{2}(2\lambda_{14}v_{0}^{2}(m_{h_{1}}^{2}-s)-4m_{h_{1}}^{4}+s^{2})}{2m_{h_{1}}^{4}-3m_{h_{1}}^{2}s+s^{2}}\log(\frac{2m_{h_{1}}^{2}+s(\sqrt{\frac{(s-4m_{h_{1}}^{2})(s-4m_{1}^{2})}{s^{2}}}-1)}{2m_{h_{1}}^{2}-s(\sqrt{\frac{(s-4m_{1}^{2})(s-4m_{h_{1}}^{2})}{s^{2}}}+1)}))$ (45)
$\sigma_{S_{1}S_{1}\to h_{2}h_{2}}$ $=\frac{\lambda_{13}^{2}}{8\pi s(s-4m_{1}^{2})}(\sqrt{(s-4m_{h_{2}}^{2})(s-4m_{1}^{2})}(\frac{8\lambda_{13}^{2}v_{1}^{4}}{m_{h_{2}}^{4}-4m_{h_{2}}^{2}m_{1}^{2}+m_{1}^{2}s}+\frac{(2m_{h_{2}}^{2}+s)^{2}}{(m_{h_{2}}^{2}-s)^{2}})$
$+$ $\frac{8\lambda_{13}v_{1}^{2}(2\lambda_{13}v_{1}^{2}(m_{h_{2}}^{2}-s)-4m_{h_{2}}^{4}+s^{2})}{2m_{h_{2}}^{4}-3m_{h_{2}}^{2}s+s^{2}}\log(\frac{2m_{h_{2}}^{2}+s(\sqrt{\frac{(s-4m_{h_{2}}^{2})(s-4m_{1}^{2})}{s^{2}}}-1)}{2m_{h_{2}}^{2}-s(\sqrt{\frac{(s-4m_{1}^{2})(s-4m_{h_{2}}^{2})}{s^{2}}}+1)}))$ (46)
$\sigma_{S_{2}S_{2}\to h_{1}h_{1}}$ $=\frac{\lambda_{24}^{2}}{8\pi s(s-4m_{2}^{2})}(\sqrt{(s-4m_{h_{1}}^{2})(s-4m_{2}^{2})}(\frac{8\lambda_{24}^{2}v_{0}^{4}}{m_{h_{1}}^{4}-4m_{h_{1}}^{2}m_{2}^{2}+m_{2}^{2}s}+\frac{(2m_{h_{1}}^{2}+s)^{2}}{(m_{h_{1}}^{2}-s)^{2}})$
$+$ $\frac{8\lambda_{24}v_{0}^{2}(2\lambda_{24}v_{0}^{2}(m_{h_{1}}^{2}-s)-4m_{h_{1}}^{4}+s^{2})}{2m_{h_{1}}^{4}-3m_{h_{1}}^{2}s+s^{2}}\log(\frac{2m_{h_{1}}^{2}+s(\sqrt{\frac{(s-4m_{h_{1}}^{2})(s-4m_{2}^{2})}{s^{2}}}-1)}{2m_{h_{1}}^{2}-s(\sqrt{\frac{(s-4m_{2}^{2})(s-4m_{h_{1}}^{2})}{s^{2}}}+1)}))$ (47)
$\sigma_{S_{2}S_{2}\to h_{2}h_{2}}$ $=\frac{\lambda_{23}^{2}}{8\pi s(s-4m_{2}^{2})}(\sqrt{(s-4m_{h_{2}}^{2})(s-4m_{2}^{2})}(\frac{8\lambda_{23}^{2}v_{1}^{4}}{m_{h_{2}}^{4}-4m_{h_{2}}^{2}m_{2}^{2}+m_{2}^{2}s}+\frac{(2m_{h_{2}}^{2}+s)^{2}}{(m_{h_{2}}^{2}-s)^{2}})$
$+$ $\frac{8\lambda_{23}v_{1}^{2}(2\lambda_{23}v_{1}^{2}(m_{h_{2}}^{2}-s)-4m_{h_{2}}^{4}+s^{2})}{2m_{h_{2}}^{4}-3m_{h_{2}}^{2}s+s^{2}}\log(\frac{2m_{h_{2}}^{2}+s(\sqrt{\frac{(s-4m_{h_{2}}^{2})(s-4m_{2}^{2})}{s^{2}}}-1)}{2m_{h_{2}}^{2}-s(\sqrt{\frac{(s-4m_{2}^{2})(s-4m_{h_{2}}^{2})}{s^{2}}}+1)}))$ (48)
$\sigma_{S_{1}S_{1}\to h_{1}h_{2}}$ $=\frac{4\lambda_{13}^{2}\lambda_{14}^{2}v_{0}^{2}v_{1}^{2}}{\pi s(s-4m_{1}^{2})}\frac{\log(\frac{m_{h_{1}}^{2}+s(\sqrt{\frac{(s-4m_{1}^{2})(m_{h_{1}}^{4}-2m_{h_{1}}^{2}(m_{h_{2}}^{2}+s)+(m_{h_{2}}^{2}-s)^{2})}{s^{3}}}-1)+m_{h_{2}}^{2}}{m_{h_{1}}^{2}-s(\sqrt{\frac{(s-4m_{1}^{2})(m_{h_{1}}^{4}-2m_{h_{1}}^{2}(m_{h_{2}}^{2}+s)+(m_{h_{2}}^{2}-s)^{2})}{s^{3}}}+1)+m_{h_{2}}^{2}})}{m_{h_{1}}^{2}+m_{h_{2}}^{2}-s}$
$+$ $\frac{s\sqrt{\frac{(s-4m_{1}^{2})(m_{h_{1}}^{4}-2m_{h_{1}}^{2}(m_{h_{2}}^{2}+s)+(m_{h_{2}}^{2}-s)^{2})}{s}}}{2(-2m_{1}^{2}s(m_{h_{1}}^{2}+m_{h_{2}}^{2})+m_{1}^{2}(m_{h_{1}}^{2}-m_{h_{2}}^{2})^{2}+m_{h_{1}}^{2}m_{h_{2}}^{2}s+m_{1}^{2}s^{2})}$ (49)
$\sigma_{S_{2}S_{2}\to h_{1}h_{2}}$ $=\frac{4\lambda_{23}^{2}\lambda_{24}^{2}v_{0}^{2}v_{1}^{2}}{\pi s(s-4m_{2}^{2})}\frac{\log(\frac{m_{h_{1}}^{2}+s(\sqrt{\frac{(s-4m_{2}^{2})(m_{h_{1}}^{4}-2m_{h_{1}}^{2}(m_{h_{2}}^{2}+s)+(m_{h_{2}}^{2}-s)^{2})}{s^{3}}}-1)+m_{h_{2}}^{2}}{m_{h_{1}}^{2}-s(\sqrt{\frac{(s-4m_{2}^{2})(m_{h_{1}}^{4}-2m_{h_{1}}^{2}(m_{h_{2}}^{2}+s)+(m_{h_{2}}^{2}-s)^{2})}{s^{3}}}+1)+m_{h_{2}}^{2}})}{m_{h_{1}}^{2}+m_{h_{2}}^{2}-s}$
$+$ $\frac{s\sqrt{\frac{(s-4m_{2}^{2})(m_{h_{1}}^{4}-2m_{h_{1}}^{2}(m_{h_{2}}^{2}+s)+(m_{h_{2}}^{2}-s)^{2})}{s}}}{2(-2m_{2}^{2}s(m_{h_{1}}^{2}+m_{h_{2}}^{2})+m_{2}^{2}(m_{h_{1}}^{2}-m_{h_{2}}^{2})^{2}+m_{h_{1}}^{2}m_{h_{2}}^{2}s+m_{2}^{2}s^{2})}$ (50)
According to the expressions of the cross section, in the limit of $\sin\theta\to 0$, the signs of $\lambda_{13}$,$\lambda_{23}$,$\lambda_{14}$ and $\lambda_{24}$ can make difference on the terms including the odd orders of these couplings so that contribute to the cross section of $S_{i}S_{i}\to h_{i}h_{i}(i=1,2)$. Note that for the non-zero $\sin\theta$, the sign of $\lambda_{13}$,$\lambda_{23}$,$\lambda_{14}$ and $\lambda_{24}$ can make also difference in the cross section by the combined contributions of these couplings such as $\lambda_{13}\lambda_{24}$.